OpenStudy (anonymous):

What is the value of the discriminant of 2x2 + 3x – 14 = 0?

OpenStudy (anonymous):

The discriminant of a quadratic is going to be b^2-4ac where the quadratic is noted as ax^2+bx+x=0

OpenStudy (anonymous):

correct that previous message; It should be ax^2+bx+c = 0

OpenStudy (anonymous):

So, in this case, a=2, b=3, and c=-14. Therefore, the discriminant is 3^2-4*2*(-14) = 9+104 = 113

OpenStudy (anonymous):

The important thing to note is that if the discriminant is negative, then there is no real solution to the quadratic. This can be inferred from the quadratic formula where x = (-b +/- sqrt(b^2-4ac))/(2a)